# MoRe: Class Patch Attention Needs Regularization for Weakly Supervised Semantic Segmentation

### v2

@@ Introduction
Weakly supervised semantic segmentation (WSSS) aims to achieve pixel-wise predictions with cheap annotations, such as points (Bearman et al. 2016), scribbles (Lin et al. 2016; Vernaza and Chandraker 2017), bounding boxes (Dai, He, and Sun 2015; Lee et al. 2021b), and image-level labels (Pinheiro and Collobert 2015; Yang et al. 2024a). It significantly reduces the annotation cost of fully supervised segmentation methods (Tang et al. 2023; Wang et al. 2022a; Xu et al. 2024b) and has attracted increasing attention in recent years. Among all annotation forms, the image-level label is the most accessible while challenging, as it contains the least semantic information to localize objects. In this work, we focus on the WSSS paradigm with only image-level labels.
@@
However, as demonstrated in (Sun et al. 2024; Darcet et al. 2023), in order to capture long-range dependencies, ViT inclines to aggregate global semantics in low-information patches. It causes unrelated patches to be frequently correlated with class tokens during attention, leading to numerous false activations of artifact patch tokens. Previous LAM-based WSSS methods barely regularize such attention. Consequently, this issue commonly exists in ViTs (such as ViT on ImageNet (Deng et al. 2009) or DeiT (Touvron et al. 2021)) and severally impairs the quality of LAM, as shown in Figure 1 (a, b). Importantly, unlike the notorious over-smoothness issue (Wang et al. 2022b) of ViT that stems from the high similarity among patch tokens, the reported artifact issue is derived from the improper correlation between class-patch tokens. Although recent WSSS works have proposed to solve the over-smoothness (Ru et al. 2023; He et al. 2023), both LAM-based and CAM-based methods particularly overlook the artifact issue from improper class-patch attention, leaving this problem unaddressed.
@@
Graph neural networks (GNN) aim to model the representation with graph elements (e.g., nodes and edges) and approximation inference, which highlights the relation within the entity topology (Wu et al. 2020). With its ability to capture complex interactions among nodes, GNN has attracted widespread attention across various fields, such as 3D pose estimation (Cai et al. 2019) and recommendation (Fan et al. 2019). For WSSS, most GNN-related works are based on convolutional neural networks (CNN). GraphNet (Pu et al. 2018) and A2GNN (Zhang et al. 2021) are proposed to construct graph structures based on superpixel or affinity under the supervision of bounding boxes. GSM (Li et al. 2021) employs GNN to capture semantic dependencies among images belonging to the same class. Although impressive, they typically rely on pre-defined relations to construct the CNN-based graph and ignore the directed interaction among entities. In this work, we build our graph on ViT and dynamically model the graph as a directed topology, which fully mines the mutual information among class-patch tokens.

### v3

@@ Introduction
Weakly supervised semantic segmentation (WSSS) aims to achieve pixel-wise predictions with cheap annotations, such as points (Bearman et al. 2016), scribbles (Lin et al. 2016; Vernaza and Chandraker 2017), bounding boxes (Dai, He, and Sun 2015; Lee et al. 2021b), and image-level labels (Pinheiro and Collobert 2015; Yang et al. 2024a). It significantly reduces the annotation cost of fully supervised segmentation methods (Tang et al. 2023; Xu et al. 2024b) and has attracted increasing attention in recent years. Among all annotation forms, the image-level label is the most accessible while challenging, as it contains the least semantic information to localize objects. In this work, we focus on the WSSS paradigm with only image-level labels.
@@
However, as demonstrated in (Sun et al. 2024; Darcet et al. 2023), in order to capture long-range dependencies, ViT inclines to aggregate global semantics in low-information patches. It causes unrelated patches to be frequently correlated with class tokens during attention, leading to numerous false activations of artifact patch tokens. Previous LAM-based WSSS methods barely regularize such attention. Consequently, this issue commonly exists in ViTs (such as ViT on ImageNet (Deng et al. 2009) or DeiT (Touvron et al. 2021)) and severally impairs the quality of LAM, as shown in Figure 1 (a, b). Importantly, unlike the notorious over-smoothness issue (Wang et al. 2022) of ViT that stems from the high similarity among patch tokens, the reported artifact issue is derived from the improper correlation between class-patch tokens. Although recent WSSS works have proposed to solve the over-smoothness (Ru et al. 2023; He et al. 2023), both LAM-based and CAM-based methods particularly overlook the artifact issue from improper class-patch attention, leaving this problem unaddressed.
@@
Graph neural networks (GNN) aim to model the representation with graph elements (e.g., nodes and edges) and approximation inference (Wu et al. 2020). With its ability to capture complex interactions within the entity topology, GNN has attracted widespread attention across various fields, such as 3D pose estimation (Mehraban et al. 2024) and medical image analysis (Zhou et al. 2019). For WSSS, most GNN-related works are based on convolutional neural networks (CNN). GraphNet (Pu et al. 2018) and A2GNN (Zhang et al. 2021) are proposed to construct graph structures based on superpixel or affinity under the supervision of bounding boxes. GSM (Li et al. 2021) employs GNN to capture semantic dependencies among images within the same class. Although impressive, they typically rely on pre-defined relations to construct the CNN-based graph and ignore the directed interaction among entities. In this work, we build our graph on ViT following (Luo, Thost, and Shi 2024; Li et al. 2024), while uniquely highlighting a directed topology between class-patch tokens to mine their mutual information.